# Optical design and sensitivity optimization of Cryogenic sub-Hz cROss torsion bar detector with quantum NOn-demolition Speed meter (CHRONOS)

### v1

@@ IV.2 Control ports and control scheme
  • POP-B (far field): Located downstream with a Gouy-phase shift of $45^{\circ}$-$90^{\circ}$ relative to POP-A, providing high sensitivity to the differential Sagnac mode (DSAG).
Both POP-A and POP-B are demodulated at $f_{2}=58.980\penalty 10000\ \mathrm{MHz}$. The Gouy-phase difference enables clear separation of the CSAG, PRCL and DSAG in detection.
Antisymmetric port (AS).
The AS port corresponds to the output (dark) port of the Sagnac interferometer. When the interferometer is on resonance, the carrier light ideally remains dark at this port (dark fringe). The demodulated signal at $f_{2}=58.980\penalty 10000\ \mathrm{MHz}$ provides the main error signal for the SRC length (SRCL) Evans et al. (2015). Meanwhile, the DC component is read out by BHD, where the velocity signal contained in the DSAG is extracted as the scientific output Fricke et al. (2012); Vahlbruch et al. (2010). The homodyne angle is set near the phase quadrature, and the in-phase component is electronically blended at low frequencies to improve control stability Danilishin et al. (2019).
Comparison with a Fabry-Perot Michelson interferometer.
The DSAG corresponds to the DARM degree of freedom in a Michelson interferometer, but in the CHRONOS configuration it represents a velocity-type observable that measures the difference in propagation speed between the CW and CCW beams rather than their positional displacement Chen (2003). Similarly, the CSAG corresponds to the Michelson CARM, representing the common path length coupled to the laser frequency and PRC resonance.
Table 1: Major detection ports and their corresponding control degrees of freedom (DOF) in the CHRONOS dual-recycled Sagnac interferometer. The top row, IMC-INI, represents the initial-alignment and frequency-reference stage in the IMC, where the length of the IMC cavity (IMCL) is maintained by a 44.235 MHz PDH signal. The subsequent ports—REFL, POP, and AS—are used to sense and control the common and differential DOFs of the main dual-recycled Sagnac interferometer (PRCL, CSAG, DSAG, and SRCL). All RF ports use phase-coherent sidebands at $f_{1}=29.500\penalty 10000\ \mathrm{MHz}$ and $f_{2}=58.980\penalty 10000\ \mathrm{MHz}$, corresponding respectively to the resonance conditions of the PRC and the SRC. The final port, AS-BHD, is a DC readout channel that extracts the velocity component of the DSAG as the speed-meter signal—the scientific output of the interferometer.
@@ V Mode-matching design
The eigenmodes of an interferometer are determined using the ABCD-matrix formalism for Gaussian beams Kogelnik and Li (1966). The complex beam parameter $q$ is defined as
@@
where beam radius $w(z)$, the radius of wavefront curvature $R(z)$, and the laser wavelength $\lambda$ are defined. An optical system is represented by the ABCD matrix
@@
For a resonant cavity, the stable eigenmode is obtained by solving $q=(Aq+B)/(Cq+D)$ for the round-trip matrix $M_{\mathrm{rt}}$. From this, the beam waist $w_{0}$, Rayleigh range $z_{R}=\pi w_{0}^{2}/\lambda$, and Gouy phase $\zeta(z)=\arctan(z/z_{R})$ are derived.
@@
where perfect overlap gives $\eta=1$. Any mismatch in waist size, curvature, or alignment decreases the coupled circulating power and increases the prompt reflection at the input port.
@@
For two coaxial Gaussian beams, the coupling efficiency can be expressed exactly in terms of the waist radii $w_{i}$ and wavefront curvature radii $R_{i}$ ($i=1,2$) Bayer-Helms (1984):
@@ V.1 Requirement conditions
In the mode-matching design of the CHRONOS triangular ring cavity, the internal optical loss $L_{\mathrm{int}}$ is sufficiently small—below $10^{-5}$ (10 ppm)—thanks to high-quality dielectric coatings. Therefore, the dominant factors that determine the optical performance are the external coupling efficiency $\eta$ and the coupling symmetry between the CW and CCW propagation paths.
@@
This quantity directly governs the balance of radiation pressure between the CW and CCW beams and determines the stability of the quantum nondemolition (QND) response in the speed-meter interferometer Chen (2003).
Derivation of the asymmetry condition.
The radiation-pressure noise appearing in the output of a speed-meter interferometer can be approximated as Chen (2003)
@@
Practical mode-matching efficiency.
When the internal loss is sufficiently low, the conventional criterion of keeping the external coupling loss $(1-\eta)$ more than an order of magnitude smaller than the internal loss is no longer realistic. Considering alignment tolerances, wavefront aberrations, and residual higher-order mode content, the practically achievable upper limit of mode-matching efficiency during design verification is approximately $\eta\simeq 99.5\%$ Kwee et al. (2012). This value is experimentally attainable and ensures both long-term operational stability and reproducibility, while fully satisfying the asymmetry condition $|\Delta\eta/\eta|<3\times 10^{-5}$.
@@
$\eta_{\mathrm{req}}\geq 99.5\penalty 10000\ \%,\qquad\left|\frac{\Delta\eta}{\eta}\right|<3\times 10^{-5}.$
These criteria guarantee that the CW and CCW radiation pressures are precisely balanced, allowing stable high-finesse operation with preserved quantum-nondemolition (QND) sensitivity.
@@ V.2 Optimization strategy
(i) In both the sagittal and tangential planes, the cavity eigenmode parameter $q_{\mathrm{cav}}$ was obtained using ABCD-matrix analysis Kogelnik and Li (1966). From the resulting mode, the wavefront curvature $R(z)$ and beam radius $w(z)$ were calculated, and the astigmatism (the difference between sagittal and tangential planes) was quantitatively evaluated.
(ii) A parameter sweep was performed with the incidence angle on the curved mirror as a variable, searching for the condition that minimizes the incidence angle. This suppresses higher-order transverse-mode excitations and minimizes the effect of wavefront aberrations.
(iii) Geometric symmetry was enforced between the left and right Sagnac rings, imposing a constraint that the CW and CCW eigenmodes must coincide. This guarantees symmetric resonance conditions and beam parameters, ensuring a balanced interferometric response.
(iv) Within these constraints, the complex beam parameter of the external injection beam, $q_{\mathrm{in}}$, was fixed, and the mismatch at the output plane was minimized via ABCD propagation:
@@
The optimal curvature $R_{\mathrm{opt}}$ was numerically determined by maximizing the mode-matching efficiency $\eta$ as the objective function, using the standard Gaussian-beam overlap expressions Bayer-Helms (1984).
(v) For the obtained $R_{\mathrm{opt}}$, the cavity stability condition $0<g_{1}g_{2}<1$ was verified, and the round-trip Gouy phase $\psi_{\mathrm{rt}}$ and transverse-mode spacing (TMS) were computed. When these values fall within the stable region and remain nondegenerate, the optimization is considered converged Kogelnik and Li (1966).
@@ V.3 Results
In the first stage of this study, an optimization was performed to maximize the mode-matching efficiency $\eta$ of the cavity, using the radius of curvature $R$ of each curved mirror as a design variable. The optimization was conducted by determining the cavity eigenmode from ABCD-matrix analysis and by searching for the conditions in which the waist position and wavefront curvature of the injected beam are best matched to those of the cavity mode Kogelnik and Li (1966); Bayer-Helms (1984).
@@
As shown in Table 6 in Appendix A, the optimal radius of curvature $R$ for each mirror was determined. Under these conditions, at the input test-mass planes (ITML and ITMR), the mode-matching efficiency for the PR–Ring and SR–Ring coupling reached $\eta\sim 99.6\%$, well exceeding the practical requirement of $\eta_{\mathrm{req}}\geq 99.5\%$ Kwee et al. (2012).
The differential beam parameters $\Delta q$, wavefront-curvature difference $\Delta R$, beam-radius difference $\Delta w$, and corresponding efficiencies $\eta$ obtained at each coupling point are summarized in Table 2. In all cases, $|\Delta q|\lesssim 9\times 10^{-2}\,\mathrm{m}$ and $|\Delta w|\lesssim 3\times 10^{-5}\,\mathrm{m}$, indicating that the coupling loss remains well below the internal-loss budget.
Table 2: Optimized mode-matching results between the PR/SR cavities and the ring cavities at the ITML and ITMR planes. Listed are the real and imaginary parts of the differential beam parameter ($\Delta q_{\mathrm{re}}$, $\Delta q_{\mathrm{im}}$), its magnitude $|\Delta q|$, differences in wavefront curvature ($\Delta R$), beam radius ($\Delta w$), and the resulting mode-matching efficiency $\eta$ after optimization. All efficiencies exceed the target threshold of $\eta\geq 0.995$. Positive $\Delta R$ and $\Delta w$ indicate a larger radius of curvature and beam size in the PR/SR modes relative to the ring mode.
@@ V.4 Stability evaluation (cavity eigenmode stability analysis)
Using the optimized curvature radii $R$ obtained from the mode-matching optimization, stability analyses were performed for the left and right ring cavities based on both ABCD-matrix calculations and Finesse3 simulations Kogelnik and Li (1966). In the evaluation, full width at half maximum (FWHM), pole frequency, and the round-trip Gouy phase $\psi_{\mathrm{rt}}$ and transverse mode spacing (TMS) for the sagittal and tangential planes were calculated Kogelnik and Li (1966).
The results are summarized in Table 4. In all cavities, the round-trip Gouy phases satisfy $0^{\circ}<\psi_{\mathrm{rt}}<180^{\circ}$, corresponding to the stable-cavity condition $0<g_{1}g_{2}<1$. In particular, the Gouy phases for the sagittal and tangential planes agree around $153^{\circ}$, confirming the formation of stable eigenmodes in both sagittal and tangential directions.
The obtained finesse, $\mathcal{F}=3.14\times 10^{4}$, and free spectral range, $\mathrm{FSR}=58.98\penalty 10000\ \mathrm{MHz}$, yield a resonance linewidth of $1.88\penalty 10000\ \mathrm{kHz}$ and a cavity pole frequency (half-width at half maximum) of $939\penalty 10000\ \mathrm{Hz}$, demonstrating that the ring cavity possesses sufficiently high optical storage time and stability Drever et al. (1983); Kwee et al. (2012).
@@
Cavity Plane $\psi_{\mathrm{rt}}$ [deg] TMS [MHz] FSR [MHz] Pole [Hz]
Left ring s 153.3 50.2 59.0 939.1
Left ring t 153.0 50.1 59.0 939.1
Right ring s 153.0 50.1 59.0 939.1
Right ring t 152.7 50.0 59.0 939.1
From these results, both the left and right ring cavities are confirmed to be stable, with the Gouy-phase difference between the s- and t-planes smaller than $0.6^{\circ}$ and the difference in transverse-mode spacing less than $0.2\penalty 10000\ \mathrm{MHz}$. This indicates excellent modal consistency between the two cavities, which contributes to suppressing higher-order mode excitation and ensuring the stable operation of the entire interferometer Kogelnik and Li (1966). Moreover, the obtained finesse and pole-frequency values satisfy the design requirements of the CHRONOS interferometer for high optical storage time and quantum-noise-limited sensitivity Kwee et al. (2012).
@@ VI Sensitivity analysis
The sensitivity of the 2.5 m CHRONOS interferometer is determined by the interplay between quantum noise and technical noise sources. In this study, the optical and mechanical parameters were set to the values listed in Table 5, which define the optimal operating point. These values are chosen to balance the contributions of shot noise, radiation-pressure noise, and coating thermal noise, optimizing the detector sensitivity around 1 Hz.
Table 5: Assumed optimal optical and mechanical parameters for the 2.5 m CHRONOS interferometer. In the CHRONOS phase convention, the detuning phase of the power-recycling cavity (PRC) is defined with respect to the $0^{\circ}$ baseline, whereas both the signal-recycling cavity (SRC) and the homodyne detection phase are referenced to the $90^{\circ}$ baseline. This convention is consistently used throughout the detuning and quadrature-rotation analysis.
@@
Ring-cavity finesse $F_{\mathrm{ring}}$ $3.14\times 10^{4}$
@@
ETM moment of inertia $I_{\mathrm{ETM}}$ $19.9\penalty 10000\ \mathrm{kg\,m^{2}}$
@@
In the first part of this section (Sec. VI.1), a complete quantum-noise model including the input-output relations and effective optomechanical coupling is introduced. In the latter part, numerical simulations using Finesse3 are presented to show the resulting sensitivity curves, focusing on the effects of PRC detuning, SRC detuning, and homodyne detection angle. Furthermore, the contribution of the ETM coating thermal noise is quantitatively evaluated, as it sets the lower sensitivity limit in the mid-frequency band Harry et al. (2002); Levin (1998). A summary of the suspension and thermal noise sources is given in the Appendix B. The intensity noise is calculated in Tanabe et. al. Tanabe et al. (2025).

### v2

@@ IV.2.3 Pick-off port (POP)
  • POP-B (far field): Located downstream with a Gouy-phase shift of $45^{\circ}$$90^{\circ}$ relative to POP-A, providing high sensitivity to the differential Sagnac mode (DSAG).
Both POP-A and POP-B are demodulated at $f_{2}=58.980\penalty 10000\ \mathrm{MHz}$. The Gouy-phase difference enables clear separation of the CSAG, PRCL, and DSAG in detection.
IV.2.4 Antisymmetric port (AS)
The AS port corresponds to the output (dark) port of the Sagnac interferometer. When the interferometer is on resonance, the carrier light ideally remains dark at this port (dark fringe). The demodulated signal at $f_{2}=58.980\penalty 10000\ \mathrm{MHz}$ provides the primary error signal for controlling the signal-recycling cavity length (SRCL) Evans2015 .
The AS signals detected in the horizontal (ASX) and vertical (ASY) sensing channels are combined in the output beam splitter (OBS) and decomposed into common and differential modes. The differential mode corresponds to the relative angular motion of the mirrors and therefore carries the scientific signal of interest. This differential AS signal is injected into the OMC, and the final science signal, denoted as DSAG, is extracted using balanced-homodyne detection fricke2012 ; vahlbruch2010 .
IV.2.5 Comparison with a Fabry–Perot Michelson interferometer
The DSAG corresponds to the DARM degree of freedom in a Michelson interferometer, but in the CHRONOS configuration it represents an angular momentum-type observable that measures the difference in propagation speed between the CW and CCW beams rather than their positional displacement chen2003 . Similarly, the CSAG corresponds to the Michelson CARM, representing the common path length coupled to the laser frequency and PRC resonance.
Table 1: Major detection ports and their corresponding control degrees of freedom in the CHRONOS dual-recycled Sagnac interferometer. The top row, IMC-INI, represents the initial-alignment and frequency-reference stage in the IMC, where the IMCL is stabilized using a 44.235 MHz Pound–Drever–Hall (PDH) signal. The subsequent ports—REFL, POP, and AS—are used to sense and control the common and differential DOFs of the main dual-recycled Sagnac interferometer (PRCL, CSAG, DSAG, and SRCL). All RF ports employ phase-coherent modulation sidebands at $f_{1}=29.500\penalty 10000\ \mathrm{MHz}$ and $f_{2}=58.980\penalty 10000\ \mathrm{MHz}$, which correspond to the resonance conditions of the PRC and the SRC, respectively. The final port, AS-BHD, is a DC readout channel that extracts the velocity component of the DSAG as the speed-meter signal, which constitutes the scientific output of the interferometer. The physical locations of the individual detection and control ports are illustrated in Fig. 2.
@@ V Mode-matching design
The eigenmodes of an interferometer are determined using the ABCD-matrix formalism for Gaussian beams Kogelnik1966 . The complex beam parameter $q$ is defined as
@@
where the beam radius $w(z)$, the radius of wavefront curvature $R(z)$, and the laser wavelength $\lambda$ are defined.
An optical system is represented by the ABCD matrix
@@
For a resonant cavity, the stable eigenmode is obtained by solving the round-trip matrix $M_{\mathrm{rt}}=\prod M$. From this, the beam waist $w_{0}$, Rayleigh range $z_{R}=\pi w_{0}^{2}/\lambda$, and Gouy phase $\zeta(z)=\arctan(z/z_{R})$ are derived.
The input beam and the cavity eigenmode are described by fundamental Gaussian beams. The complex electric fields are written as
$E_{\mathrm{in}}(x,y,z)=E_{0,\mathrm{in}}\frac{w_{0,\mathrm{in}}}{w_{\mathrm{in}}(z)}\exp\left[-\frac{x^{2}+y^{2}}{w_{\mathrm{in}}^{2}(z)}\right]\exp\left[-ikz-\frac{ik(x^{2}+y^{2})}{2R_{\mathrm{in}}(z)}+i\zeta_{\mathrm{in}}(z)\right],$ (13)
$E_{\mathrm{cav}}(x,y,z)=E_{0,\mathrm{cav}}\frac{w_{0,\mathrm{cav}}}{w_{\mathrm{cav}}(z)}\exp\left[-\frac{x^{2}+y^{2}}{w_{\mathrm{cav}}^{2}(z)}\right]\exp\left[-ikz-\frac{ik(x^{2}+y^{2})}{2R_{\mathrm{cav}}(z)}+i\zeta_{\mathrm{cav}}(z)\right],$ (14)
where $k=2\pi/\lambda$ is the wave number, $w(z)$ is the beam radius, $R(z)$ is the wavefront radius of curvature, and $\zeta(z)$ is the Gouy phase. These parameters are related through the complex beam parameter $q(z)$ as
$\frac{1}{q(z)}=\frac{1}{R(z)}-i\frac{\lambda}{\pi w^{2}(z)}.$ (15)
@@
where $E_{\mathrm{in}}$ represents the injected spatial mode and $E_{\mathrm{cav}}$ denotes the fundamental eigenmode of the resonant cavity. Perfect overlap gives $\eta=1$. Any mismatch in waist size, wavefront curvature, or alignment decreases the coupled circulating power and increases the prompt reflection at the input port.
@@
For two coaxial Gaussian beams, the coupling efficiency can be expressed exactly in terms of the waist radii $w_{i}$ and wavefront curvature radii $R_{i}$ ($i=1,2$) Bayer1984 :
@@ V.1 Requirement conditions
In the mode-matching design of the CHRONOS triangular ring cavity, the internal optical loss $L_{\mathrm{int}}$ is sufficiently small—below $10^{-5}$ (10 ppm)— thanks to high-quality dielectric coatings. Therefore, the dominant factors that determine the optical performance for interferometer control are the external coupling efficiency $\eta$ and the coupling symmetry between the CW and CCW propagation paths.
Here it should be emphasized that the internal optical loss discussed above has only a minor impact on the control performance. Its impact on the sensitivity is a separate issue. The sensitivity is affected through the achievable cavity finesse, which depends on the total optical loss. This aspect is discussed in detail in Sec. VI.1.
@@
This quantity directly governs the balance of radiation pressure between the CW and CCW beams and determines the stability of the quantum nondemolition (QND) response in the speed-meter interferometer chen2003 .
Derivation of the asymmetry condition
The radiation-pressure noise appearing in the output of a speed-meter interferometer can be approximated as chen2003
@@
Practical mode-matching efficiency
When the internal loss is sufficiently low, the conventional criterion of keeping the external coupling loss $(1-\eta)$ more than an order of magnitude smaller than the internal loss is no longer realistic. Considering alignment tolerances, wavefront aberrations, and residual higher-order mode content, the practically achievable upper limit of mode-matching efficiency during design verification is approximately $\eta\simeq 99.5\%$ Kwee2012 . This value is experimentally attainable and ensures both long-term operational stability and reproducibility, while fully satisfying the asymmetry condition $|\Delta\eta/\eta|<3\times 10^{-5}$.
@@
$\eta\geq 99.5\penalty 10000\ \%,\qquad\left|\frac{\Delta\eta}{\eta}\right|<3\times 10^{-5}.$
These criteria guarantee that the CW and CCW radiation pressures are precisely balanced, allowing stable high-finesse operation with preserved QND sensitivity.
@@ V.2 Optimization strategy
  1. (i)
    In both the sagittal and tangential planes, the cavity eigenmode parameter $q_{\mathrm{cav}}$ was obtained using ABCD-matrix analysis Kogelnik1966 . From the resulting mode, the wavefront curvature $R(z)$ and beam radius $w(z)$ were calculated, and the astigmatism (the difference between sagittal and tangential planes) was quantitatively evaluated.
  2. (ii)
    A parameter sweep was performed with the incidence angle on the curved mirror as a variable, searching for the condition that minimizes the incidence angle. This suppresses higher-order transverse-mode excitations and minimizes the effect of wavefront aberrations.
  3. (iii)
    Geometric symmetry was enforced between the left and right Sagnac rings, imposing a constraint that the CW and CCW eigenmodes must coincide. This guarantees symmetric resonance conditions and beam parameters, ensuring a balanced interferometric response.
  4. (iv)
    Within these constraints, the complex beam parameter of the external injection beam, $q_{\mathrm{in}}$, was fixed, and the mismatch at the output plane was minimized via ABCD propagation:
@@
    The optimal curvature $R_{\mathrm{opt}}$ was numerically determined by maximizing the mode-matching efficiency $\eta$ as the objective function, using the standard Gaussian-beam overlap expressions Bayer1984 .
  5. (v)
    For the obtained $R_{\mathrm{opt}}$, the cavity stability condition $0<g_{1}g_{2}<1$ was verified, and the round-trip Gouy phase $\psi_{\mathrm{rt}}$ and transverse-mode spacing (TMS) were computed. When these values fall within the stable region and remain nondegenerate, the optimization is considered converged Kogelnik1966 .
@@ V.3 Results
In the first stage of this study, an optimization was performed to maximize the mode-matching efficiency $\eta$ of the cavity, using the radius of curvature $R$ of each curved mirror as a design variable. The optimization was conducted by determining the cavity eigenmode from ABCD-matrix analysis and by searching for the conditions in which the waist position and wavefront curvature of the injected beam are best matched to those of the cavity mode Kogelnik1966 ; Bayer1984 .
@@
As shown in Table 6 in Appendix A, the optimal radius of curvature $R$ for each mirror was determined. Under these conditions, at the input test-mass planes (ITML and ITMR), the mode-matching efficiency for the PR–Ring and SR–Ring coupling reached $\eta\sim 99.5\%$, well exceeding the practical requirement of $\eta\geq 99.5\%$ as shown in Table 2 Kwee2012 .
The differential beam parameters $\Delta q$, wavefront-curvature difference $\Delta R$, beam-radius difference $\Delta w$, and corresponding efficiencies $\eta$ obtained at each coupling point are summarized in Table 2. In all cases, $|\Delta q|\ll 9\times 10^{-2}\,\mathrm{m}$ and $|\Delta w|\ll 3\times 10^{-5}\,\mathrm{m}$, indicating that the coupling loss remains well below the internal-loss budget.
Table 2: Optimized mode-matching results between the PR/SR cavities and the ring cavities at the ITML and ITMR planes. Listed are the real and imaginary parts of the differential beam parameter ($\Delta q_{\mathrm{re}}$, $\Delta q_{\mathrm{im}}$), its magnitude $|\Delta q|$, differences in wavefront curvature ($\Delta R$), beam radius ($\Delta w$), and the resulting mode-matching efficiency $\eta$ after optimization. All efficiencies exceed the target threshold of $\eta\geq 99.5\%$. Positive $\Delta R$ and $\Delta w$ indicate a larger radius of curvature and beam size in the PR/SR modes relative to the ring mode.
@@ V.4 Stability evaluation (cavity eigenmode stability analysis)
Using the optimized curvature radii $R$ obtained from the mode-matching optimization, stability analyses were performed for the left and right ring cavities based on both ABCD-matrix calculations and Finesse3 simulations Kogelnik1966 . In the evaluation, full width at half maximum (FWHM), pole frequency, and the round-trip Gouy phase $\psi$ and transverse mode spacing (TMS) for the sagittal and tangential planes were calculated Kogelnik1966 .
The results are summarized in Table 4. In all cavities, the round-trip Gouy phases satisfy $0^{\circ}<\psi<180^{\circ}$, corresponding to the stable-cavity condition $0<g_{1}g_{2}<1$. In particular, the Gouy phases for the sagittal and tangential planes agree around $153^{\circ}$, confirming the formation of stable eigenmodes in both sagittal and tangential directions.
The obtained finesse, $\mathcal{F}=3.14\times 10^{4}$, and free spectral range, $\mathrm{FSR}=58.98\penalty 10000\ \mathrm{MHz}$, yield a resonance linewidth of $1.88\penalty 10000\ \mathrm{kHz}$ and a cavity pole frequency (half-width at half maximum) of $939\penalty 10000\ \mathrm{Hz}$, demonstrating that the ring cavity possesses sufficiently high optical storage time and stability Drever1983 ; Kwee2012 . From these results, both the left and right ring cavities are confirmed to be stable, with the Gouy-phase difference between the s- and t-planes smaller than $0.6^{\circ}$ and the difference in transverse-mode spacing less than $0.2\penalty 10000\ \mathrm{MHz}$. This indicates excellent modal consistency between the two cavities, which contributes to suppressing higher-order mode excitation and ensuring the stable operation of the entire interferometer Kogelnik1966 . Moreover, the obtained finesse and pole-frequency values satisfy the design requirements of the CHRONOS interferometer for high optical storage time and quantum-noise-limited sensitivity Kwee2012 .
@@
Cavity Plane $\psi$ [deg] TMS [MHz] FSR [MHz] Pole [Hz]
Left ring s 153.3 50.2 59.0 939
Left ring t 153.0 50.1 59.0 939
Right ring s 153.0 50.1 59.0 939
Right ring t 152.7 50.0 59.0 939
@@ VI Sensitivity analysis
The sensitivity of the $2.5\penalty 10000\ \mathrm{m}$ CHRONOS interferometer is determined by the interplay between quantum noise and technical noise sources. In this study, the optical and mechanical parameters were set to the values listed in Table 5, which define the optimal operating point. These values are chosen to balance the contributions of shot noise, radiation-pressure noise, coating thermal noise, bar thermal noise, and Newtonian noise, thereby optimizing the detector sensitivity around $1\penalty 10000\ \mathrm{Hz}$.
Table 5: Assumed optimal optical and mechanical parameters for the 2.5 m CHRONOS interferometer. In the CHRONOS phase convention, the detuning phase of the PRC is defined with respect to the $0^{\circ}$ baseline, whereas both the SRC and the homodyne detection phase are referenced to the $90^{\circ}$ baseline. This convention is consistently used throughout the detuning and quadrature-rotation analysis.
@@
Ring-cavity finesse $\mathcal{F}$ $3.14\times 10^{4}$
@@
ETM moment of inertia II $19.9\penalty 10000\ \mathrm{kg\,m^{2}}$
@@
In the first part of this section (Sec. VI.1), a complete quantum-noise model including the input-output relations and effective optomechanical coupling is introduced. In the latter part, numerical simulations using Finesse3 are presented to show the resulting sensitivity curves, focusing on the effects of ETM reflectance, PRC detuning, SRC detuning, and homodyne detection angle. A summary of the noise sources is given in Appendix B. The intensity noise is calculated in Tanabe et al. Tanabe:inprep .